# Machine-Readable Ads: Accessibility and Behavioral Patterns of AI Web Agents interacting with Online Advertisements

### v1

@@ 5.3 Limitations and Future Work
Model transparency.
Commercial stacks expose little beyond an API, making it impossible to tell whether a failure arises from pre-training data, the system prompt, or the safety layer. Re-implementing a framework like Browser Use on self-trained, open-weights backbones would change that. By running identical browsing tasks across multiple backbones while independently toggling data, prompts, and policy modules, researchers could (i) isolate which component triggers each trust-boundary breach and (ii) systematically classify entire architectures by their trustworthiness—quantifying robustness, privacy leakage, and safety-layer controllability under uniform conditions.
Prompting catalogues.
The field lacks a public, validated library of task–prompt–expected-outcome triples. A standardized catalogue would sharpen empirical comparisons, much as GLUE [^15] https://gluebenchmark.com/ and SuperGLUE [^16] https://super.gluebenchmark.com/ did for natural language processing. Although our task set was informed by face validity and domain expertise, we acknowledge that it may not fully capture the most common, real-world interaction patterns on contemporary news sites. We urge the community to curate and version these benchmarks. Doing so demands the collection of authentic, real-world interaction logs that (i) confirm each task arises naturally in practice and (ii) demonstrate that the accompanying prompt wording accurately reflects language users already employ.
Site-layout diversity.
Modern e-commerce sites vary widely—from Western grid layouts to mobile-first, vertically stacked designs common in East Asia. A systematic study of agent performance across these design archetypes would clarify which UI conventions hinder or help autonomous browsing.

### v2

@@ 5.3 Limitations and Future Work
This study set out to examine how contemporary multimodal AI web agents perceive and act within an ad-rich news environment, and to derive practical design implications for making online advertisements machine-readable without degrading the human user experience or eroding user trust. While the results are informative and, we believe, actionable, they should be interpreted in light of several limitations that stem from the experimental substrate, the nature of current commercial models, and the pragmatic scope of our measurements.
A first limitation concerns ecological validity. Our evaluation was performed on a high-fidelity clone of a production news site rather than on the live domain. This decision afforded tight control over structure, flows, and the repeatability of interaction sequences. The clone reproduces the DOM, layout, and representative creatives with care, but real-world timing, sequencing, and personalization effects remain beyond our present scope. Consequently, our findings should be interpreted as conservative lower-bound indicators of agent-ad interaction and trust-impact patterns likely to occur in the wild, rather than a comprehensive characterization of all behaviors across live ad-tech ecosystems.
A second limitation follows from model opacity and drift. The commercial stacks we exercised (e.g., GPT-4o, Claude 3.7 Sonnet, Gemini 2.0 Flash, and Operator) expose minimal internals and evolve continuously. We cannot causally disaggregate the relative influence of pretraining data, system prompts, tool-use policies, or safety layers on any specific behavior, nor can we exclude version drift over time beyond the configurations we document. This opacity is typical at present but it constrains mechanistic explanations: we can say what happened and under which conditions, but not always why at the component level.
Third, there is an instrumentation boundary between DOM-centric and pixel-centric paradigms. Our Browser-Use based runs reason over the DOM and therefore cannot directly act on native browser UI layers (e.g., permission dialogs), while the vision-first Operator bears different constraints, including higher latency, cost profiles, and explicit confirmation pauses for sensitive actions. As a result, cross-paradigm comparisons are indicative rather than strictly like-for-like. Differences we observe may sometimes reflect architectural affordances as much as model proclivities.
Fourth, our task set and prompt wording are intentionally realistic yet necessarily scoped. We centered a single news context with tasks spanning subscriptions, coupon discovery, article access, and consent handling. Other verticals (e-commerce, banking, social media), devices (especially mobile-first experiences), locales and writing styles may elicit distinct behaviors. Likewise, our instructions were authored rather than mined from user logs; while they mirror naturalistic goals, they cannot cover the full diversity of how end users phrase requests.
Fifth, our measurements emphasize concrete, observable patterns over composite indices. We deliberately refrained from introducing new trust or compliance scores and did not undertake user studies. Instead we report the salient regularities we repeatedly observed: shallow scrolling and satisficing, strong reliance on semantic overlays for clickability, and a willingness to proceed with purchases when progress is explicitly gated by paywalls. These qualitative characterizations are, we contend, valuable for design and risk assessment, but they do not substitute for agent-specific standardized policy-aware metrics.
Sixth, although our trials demonstrated that agent attention can effectively be steered using non-visible cues, our experiments tested only a small subset of possible hidden-text strategies. We also tested using only a single natural language description. The specific wording and structure of these guiding descriptions could strongly influence whether an agent recognizes and selects the content element for interaction. To advance this work, we need a comprehensive comparison of how web agent attention can be guided, taking into account exactly which parts of the DOM the agents receive via the given framework or tool, through approaches like markup, off-screen text, ARIA, and visual cues. The overarching aim is to develop a widely recognized set of guidelines for crafting these informative, machine readable, cues, prompting website developers and framework creators alike to follow standardized rules that enhance how agents interact with web content.
Finally, there are ethical and procedural constraints. We used dummy payment credentials to prevent real transactions, which safeguards participants but omits real-world frictions such as identity checks, cancellations, and emotional responses to payment prompts. Our study prioritizes control and clear observation over breadth, and future work should extend this through targeted ablations and policy-aware scenarios that make agent trust and machine-readability both actionable and measurable across models, sites, and time.